Find the probability of drawing a heart or a black card from a standard deck of cards.
: Find the probability of drawing a heart or a black card from a standard deck of cards.

Answer by checkley77(3397)

(13+26)/52=39/52=.75=75% probability of drawing a heart (13/52)or a black (26/52) card.